Please help me solve this
calculate the slope of the line passing through each pair of points. Then tell whether the line rises, falls, is horizontal, or is vertical?
(3,5) and (9,15) slope=________,_________

Answer by Alan3354(69443)

Please help me solve this
calculate the slope of the line passing through each pair of points. Then tell whether the line rises, falls, is horizontal, or is vertical?
(3,5) and (9,15) slope=________,_________
--------------------
The slope, m, is the diff in y/diff in x.
m = (15-5)/(9-3)
m = 10/6 = 5/3
The slope is positive, so it "rises" from left to right (as x increases).